OpenStudy (anonymous):

i want the final solution 4: express cos 7 theta in terms of cos theta.. can u help?

OpenStudy (zehanz):

using x instead of theta - makes typing easier ;) cos7x= cos(3x+4x)= cos3xcos4x-sin3xsin4x (1) Now I will first try to do away with cos3x: (using cos2x=2cos²x-1, sin2x=2sinxcosx and sin²x+cos²x=1) cos3x= cos(x+2x)= cosxcos2x-sinxsin2x= cosx(2cos²x-1)-sinx2sinxcosx= cosx(2cos²x-1-2sin²x)= cosx(2cos²x-1-2(1-cos²x))= cosx(4cos²x-3). Now there is only cosx in this part of formula (1). You see this is already a lot of work. Now all you have to do is go on and rewrite the other components of (1), using the same techniques!